# WiFo-M2: Plug-and-Play Multi-Modal Sensing via Foundation Model to Empower Wireless Communications

### v1

@@ V-E3 V1 Dataset
The V1 dataset provides only images and CSI data. Hence, we focus on the performance of WiFo-M2-Img. In the BP task, the WiFo-M2-Img-based scheme achieves top‑1/3/5 accuracy of $67.4\%$, $91.2\%$, and $98.1\%$, outperforming Vision‑BP ($65.5\%$, $82.4\%$, $92.3\%$) and closely matching the performance of the MM‑BP scheme ($71.3\%$, $87.4\%$, $99.2\%$) which uses only its image branch. Table VI presents the enhancement effect of WiFo-M2-Img on other baselines, demonstrating consistent performance improvements across all baselines, with the exception of the Transformer-based CP scheme, which failed to converge. In the same way, the weights derived from the ContraSoM strategy play a significant role.
TABLE VII: Additional fine-tunable parameters introduced by applying WiFo-M2 to different transceiver modules.
Transceiver Modules CE: CENN CE: FCDAMP CI: LPCCNet BP
@@
Transceiver Modules CI: SR-CI CP: WiFo CP: Transformer
@@ V-F Computational and Deployment Efficiency
To assess the practical deployment overhead of WiFo-M2, we evaluate its model complexity and inference latency. All experiments are conducted on a server equipped with dual 14-core Intel Xeon E5-2680 v4 CPUs and an NVIDIA GeForce RTX 4090 GPU, processing inputs with a batch size of $1$ to simulate real-time operation. WiFo-M2-LiDAR comprises only 0.070.07M parameters with an inference latency of 3.313.31ms, while WiFo-M2-Img contains 9.629.62M parameters with an inference latency of 6.046.04ms. When multiple modalities are available, the total inference can be executed in (at most) the sum of their latencies, i.e., 9.359.35ms, when processed serially. Even with this upper bound, the system stays well within a real-time budget: at $40$ FPS a new multi-modal sample arrives every 2525ms, leaving 15.715.7ms after inference to exploit the freshly produced features. During those 9.359.35ms, the features predicted from the previous sample can still be used, which only requires the model’s output sequence to be extended to cover the 9.359.35ms inference window
As another practical metric, Table VII presents the additional fine-tunable parameters introduced when WiFo-M2 is integrated into various existing transceiver modules. The parameter overhead remains extremely low, reaching at most 0.260.26M. This is due to the effectiveness and superiority of the ContraSoM pre-training strategy, which allows the WiFo-M2 backbone to remain frozen during adaptation and only requires the lightweight adapter to be fine-tuned together with the existing schemes. This further highlights the practical deployability of WiFo-M2.
@@ VI Conclusions
This paper proposed the first universal foundation model for realizing multi-modal sensing-aided communications, WiFo-M2. To bridge the huge gap between multi-modal sensing and communications, we crafted the ContraSoM pre-training strategy, a contrastive pre-training scheme that leverages diffusion-based augmentation. Empowered by large-scale pre-training, WiFo-M2 is able to derive fine-grained OOB channel features from sparse historical sensing input, providing a transferable representation that plugs seamlessly into diverse transceiver modules. WiFo-M2 fundamentally shifts the paradigm from designing isolated, task-specific schemes to flexibly applying the universal representation to optimize existing transceiver design schemes, thereby extending the benefit of multi-modal sensing to a wider range of core transceiver design modules.

### v2

@@ V-E3 V1 Dataset
The V1 dataset provides only images and CSI data. Hence, we focus on the performance of WiFo-M2-Img. In BP action, the WiFo-M2-Img-based scheme achieves top‑1/3/5 accuracy of $67.4\%$, $91.2\%$, and $98.1\%$, outperforming Vision‑BP ($65.5\%$, $82.4\%$, $92.3\%$) and closely matching the performance of the MM‑BP scheme ($71.3\%$, $87.4\%$, $99.2\%$) which uses only its image branch. Table VI presents the enhancement effect of WiFo-M2-Img on other baselines, demonstrating consistent performance improvements across all baselines, with the exception of the Transformer-based CP scheme, which failed to converge. In the same way, the weights derived from the ContraSoM strategy play a significant role.
TABLE VII: Additional fine-tunable parameters introduced by applying WiFo-M2 to different PHY actions.
PHY Actions CE: CENN CE: FCDAMP CI: LPCCNet BP
@@
PHY Actions CI: SR-CI CP: WiFo CP: Transformer
@@ V-F Computational and Deployment Efficiency
To assess the practical deployment overhead of WiFo-M2, we evaluate its model complexity and inference latency. All experiments are conducted on a server equipped with dual 14-core Intel Xeon E5-2680 v4 CPUs and an NVIDIA GeForce RTX 4090 GPU, processing inputs with a batch size of $1$ to simulate real-time operation. WiFo-M2-LiDAR comprises only 0.070.07M parameters with an inference latency of 3.313.31ms, while WiFo-M2-Img contains 9.629.62M parameters with an inference latency of 6.046.04ms. When multiple modalities are available, the total inference can be executed in (at most) the sum of their latencies, i.e., 9.359.35ms, when processed serially. Even with this upper bound, the system stays well within a real-time budget: at $40$ fps a new multi-modal sample arrives every 2525ms, leaving 15.715.7ms after inference to exploit the freshly produced features. During those 9.359.35ms, the features predicted from the previous sample can still be used, which only requires the model’s output sequence to be extended to cover the 9.359.35ms inference window.
As another practical metric, Table VII presents the additional fine-tunable parameters introduced when WiFo-M2 is integrated into PHY actions. The parameter overhead remains extremely low, reaching at most 0.260.26M. This is due to the effectiveness and superiority of the ContraSoM pre-training strategy, which allows the WiFo-M2 backbone to remain frozen during adaptation and only requires the lightweight adapter to be fine-tuned together with the existing schemes. This further highlights the practical deployability of WiFo-M2.
@@ VI Conclusions
In this paper, we have realized a framework where intelligent embodied network entities utilize multi-modal environment sensing to comprehensively refine their PHY actions in a plug-and-play manner. Central to our approach is a powerful foundation model, WiFo-M2, which is capable of extracting fine-grained OOB channel-aware features from sparse historical sensing inputs. This is enabled by ContraSoM, a contrastive pre-training scheme that aligns environment sensing with wireless channels in a shared latent space. To enhance representation robustness, we further incorporate modality-specific data augmentation techniques that apply stochastic pixel-space transformations for images and leverage a diffusion-based feature generation framework for LiDAR point clouds. Extensive experiments on typical PHY actions show that our proposed framework boosts network performance from on multiple fronts and generalizes well across diverse scenarios.